Operation Summary for December 6, 2010 - December 8, 2010 |
Monday Dec. 6 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (185.5E10), with the Antiproton Source (Pbar) stacking (33.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI, MiniBooNE, and the Meson FTBF T-1008 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations established HEP store 8338 at 8:33 AM with an average initial luminosity (AIL) of 243.7E30. Water personnel repaired a LCW pump at MI-30 and RF personnel repaired a Booster RF station (BRF14). The networking group swapped some hardware that caused some communication issues. Operations reported other minor downtimes caused by Linac, Booster, and MI RF. During the evening shift, Operations reported spending a majority of the shift tuning the transfer line for Pbar due to BPM problems. Operators also tuned the Proton Source. Operations disabled beam to T-1008 at 6 PM. Tuesday Dec. 7 The midnight shift began with Operations monitoring store 8338 (52.64E30) and the stash (332.3E10), with Pbar stacking (17.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8338 at 1:10 AM, put the TeV into shot setup, and established store 8339 at 2:33 AM with an AIL of 327.55E30. Operations established beam to T-1008 at 4:20 AM. Operators tuned MI to improve efficiencies. During the day shift, MI personnel began their abort gap clearing kicker studies. A Linac RF station (LRF2) started causing problems near the end of shift. During the evening shift, Operations terminated store 8339 at 6:30 PM, put the TeV into shot setup, and established store 8340 at 8:45 PM with an AIL of 330.39E30. Operations disabled beam to T-1008 at 6 PM; this ended their run at MTest. Operations reported that MIRF16 developed a modulator problem and would not reset; operators turned off the station for the night. As a result NuMI intensities decreased. Operations reported that the first transfer of antiprotons after the shot was lost due to a kicker pre-fire; the kicker reset. Wednesday Dec. 8 The midnight shift began with Operations monitoring store 8340 (167.78E30) and the stash (193.8E10), with Pbar stacking (16.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators conducted MI tuning to bring NuMI back up to the desired intensity. Operations reported that LRF2 was still causing problems. Operations also reported the loss of small bits of luminosity due to two instances of TeV orbit monitor instabilities. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 12/08/2010 email Fermilab |
| Security, Privacy, Legal | |